# The Low-Temperature Phenomenology of Gap Inhomogeneity in the Cuprate Superconductors: High-Energy Granularity, Low-Energy Homogeneity, and Spectral Kinks

### v1

@@ II.3 Order Parameter Proxies
A central motivation of this paper is figuring out the extent to which the various observables that can be extracted from $dI/dV$ measurements can serve as accurate proxies for the true underlying superconducting order parameter. In much of the STS literature, the energy at which the superconducting coherence peak is found—the “spectral gap”—has long served as a convenient proxy for the superconducting gap (on the other hand, it should be noted that many other STS papers interpret the spectral gap as the pseudogap energy). For a bulk superconductor in the clean limit, this correspondence is exact. However, in the presence of disorder, this is no longer automatically true. When chemical-potential disorder is present, the correlation between the spectral gap and the underlying order parameter is not strong, and in certain circumstances becomes very weak Sulangi et al. (2021). As this paper focuses purely on order-parameter inhomogeneity, we will reexamine the correlation between the spectral gap and the order parameter in detail, and in particular focus on the role played by the inhomogeneity length scale $l$ on this correlation.
@@
We will focus on two proxies for the local superconducting order parameter. The first one is simply the aforementioned spectral gap $\Delta_{S}$ and is obtained by finding the location of the positive-energy peak in the LDOS. (The negative-energy spectra are disregarded as these are contaminated by the presence of the Van Hove singularity right below the Fermi energy.) The second one is christened the “low-energy gap” $\Delta_{L}$ and is obtained by taking the slope of the positive-energy LDOS near zero energy. In a bulk clean $d$-wave superconductor, the slope near $\omega=0$ is inversely proportional to the bulk order parameter, so one can use the former to obtain the latter. We use this relationship $d\rho/d\omega|_{\omega\approx 0^{+}}\propto 1/\Delta_{L}$ to obtain an estimate for $\Delta_{L}$ from the low-energy slope of the LDOS at each lattice site. To be more precise, because we use a finite frequency grid in evaluating the LDOS, we use the finite difference quotient to approximate the low-energy slope of the LDOS:
@@
While this estimate is exact only for a homogeneous $d$-wave superconductor, we will nevertheless use Eqs. 9 and 10 to estimate the low-energy gap $\Delta_{L}$ for each LDOS spectrum for the inhomogeneous case.
These two proxies are to be compared to the “true” superconducting order parameter $\Delta_{T}$, which we define in the following way:
@@
We also retain information about the coherence-peak height $\rho_{C}(\mathbf{r}_{i})=\rho(\mathbf{r}_{i},\omega=+\Delta_{S,i})$ as its correlations with the order-parameter proxies can potentially provide valuable clues about the nature of the order-parameter inhomogeneity in real-world cuprates. In particular, it is known that $\rho_{C}$ and $\Delta_{S}$ are negatively correlated for many cuprates studied by STS such as Bi-2212 and Bi-2201 Lang et al. (2002); McElroy et al. (2005a); Fang et al. (2006); Alldredge et al. (2008); Zou et al. (2022), so this provides an important experimental constraint on various theoretical models of order-parameter inhomogeneity.
@@ III Small Patches ($l<\xi$)
Figure 2: Plots of $\Delta_{T}$ (a), $\Delta_{S}$ (b), and $\Delta_{L}$ (c), and the absolute value of their respective Fourier transforms (d, e, and f) for patch size $l=1$. Only one disorder realization is presented. Shown here are quantities taken from the middlemost $256\times 256$ segment of the full $1024\times 256$ system.
@@
We first turn our attention to small patches characterized by $l<\xi$. To exemplify this, we show in Fig. 2 real-space and Fourier-transformed plots of $\Delta_{T}$, $\Delta_{S}$, and $\Delta_{L}$ for a $256\times 256$ subset of the system for the case where the mean order parameter is $\overline{\Delta}=0.16$ and the patch size is $l=1$. As mentioned earlier, for $\overline{\Delta}=0.16$, the coherence length is $\xi\approx 4.53$, so the patches are parametrically smaller than the coherence length. It can be seen from the real-space plot of the order parameter $\Delta_{T}$ (Fig. 2a) that it has a broad distribution ranging from $\Delta_{T}\approx 0$ to $\Delta_{T}\approx 0.3$, and that its modulations are short-ranged, with $\Delta_{T}$ on nearby sites largely uncorrelated with one another, except for rare regions where multiple patches with similar values are close to one another. The distribution of $\Delta_{T}$ is broad in that both regions with very small gaps and regions with very large gaps are present, often in close proximity with one another. In contrast, the spectral gap $\Delta_{S}$ (Fig. 2b) is comparatively narrowly distributed around $0.16$, except for a very small number of regions whose value clusters around $0.25$. The latter effect is due to the Van Hove singularity (VHS); in the clean limit with $\overline{\Delta}=0.16$, the VHS gives rise to a small shoulder at $\omega\approx 0.25$ above the positive-energy coherence peak at $\omega=0.16$. In the inhomogeneous setting, it appears that interference effects can cause this shoulder to mix with the coherence peak, leading to the formation of a peak where the shoulder should be. These clusters tend to form at large-$\Delta_{T}$ regions, but this effect appears to be rare as only a small number of large-$\Delta_{T}$ regions exhibit this effect at all.
@@
The real-space plot of $\Delta_{S}$ is characterized by weak modulations with a wavelength of around a few lattice sites. If one closely examines the $\Delta_{S}$ and $\Delta_{T}$ plots side by side, it can be seen that the two maps do not look similar to one another at all. While some large-$\Delta_{T}$ areas coincide with areas where $\Delta_{S}$ is large, the periodic modulations seen in $\Delta_{S}$ do not at all correspond to any modulations seen in $\Delta_{T}$. These periodic modulations are instead due to QPI arising from the presence of short-range disorder in the pairing (or “off-diagonal”) channel. The Fourier transform of the $\Delta_{S}$ map (Fig. 2e) confirms this: the dominant wavevectors here are precisely the scattering wavevectors responsible for QPI near the $d$-wave gap edge. This explanation is not altogether surprising. In the $l<\xi$ regime, the order-parameter inhomogeneity can be regarded as being in the pointlike limit in which the allowed scattering processes are unrestricted except for the requirement that they connect quasiparticle states for which the order parameter has the same sign. In the $l=1$ case, the modulations in the LDOS at $\omega\approx 0.16$ themselves form the modulations in the $\Delta_{S}$ maps (recall that $\Delta_{S}$ is the location of the positive-energy coherence peak). A close examination of Fig. 2e shows that the dominant scattering wavevectors giving rise to the modulations in the $\Delta_{S}$ map are the so-called sign-preserving “$\mathbf{q}$-vectors” well-known from the octet-model description of QPI: $\mathbf{q}_{1}$, $\mathbf{q}_{4}$, and $\mathbf{q}_{5}$ Nunner et al. (2006); Pereg-Barnea and Franz (2008); Sulangi et al. (2017).
@@
$\Delta_{L}$ (Fig. 2c) meanwhile can be seen to be very narrowly distributed around 0.16 as well, and does not reflect the comparatively large spatial variations in the underlying order parameter $\Delta_{T}$. Modulations can be seen here as well, but these are evidently very weak, and have a different characteristic length scale that those seen in $\Delta_{S}$. The Fourier transform of $\Delta_{L}$ (Fig. 2f) shows that these modulations arise from QPI as well, but this time at low energies near $\omega\approx 0$. The Fourier transforms (with visible QPI scattering wavevectors) of the two order-parameter proxies in Figs. 2e and 2f are in striking contrast to the Fourier transform of the true order parameter $\Delta_{T}$ (Fig. 2d), which is by and large featureless.
@@
Fig. 3b shows a two-dimensional histogram of the LDOS values across the same ensemble of spectra shown in the previous plot, with the disorder- and system-averaged LDOS shown in white. For each value of the frequency $\omega$, the distribution of the LDOS is obtained, and the density of each possible LDOS value is shown in Fig. 3b. It can be seen that the LDOS distribution does not deviate much from the average LDOS. This is especially true at low energies. As the energy is increased, one can see that the distribution becomes broader as ±$\to\pm\overline{\Delta}$, but the values remain within a narrow region surrounding the average LDOS.
More insight into the homogeneous nature of the states comes by way of plotting the LDOS itself (Fig. 4) taken from the same field of view as in Fig. 2 at four representative energies, which we take to be: $\omega=0.03$ (Fig. 4a), to represent the low-energy LDOS; $\omega=0.15$ and $\omega=0.17$ (Figs. 4b and d), which are the spectral gaps of the LDOS averaged over the first and last bins, respectively, in Fig. 3b; and $\omega=0.16$, which corresponds to the average superconducting order parameter in the system. The latter three energies are chosen to exemplify the high-energy states where inhomogeneity is strongest. The low-energy LDOS (Fig. 4a) is very featureless and does not contain any strong variations reflective of the underlying order parameter. As for the high-energy plots (Figs. 4b-d), these all are very similar to one another in that they feature modulations due to QPI. These moduluations are particularly pronounced for $\omega=0.15$ (Fig. 4b), but they appear to weaken as energy is increased from this point. These LDOS patterns are consistent with QPI due to weak, narrowly distributed order-parameter inhomogeneity as previously studied by Sulangi et al. Sulangi et al. (2017), but it is nevertheless striking that the same patterns are still present here even when the width of the order-parameter disorder distribution is parametrically much larger than in the case previously studied. This may be attributed to the fact that even with the relatively large width of the order-parameter inhomogeneity distribution we have used (on the order of the mean superconducting order parameter itself), it is still weak on the overall scale set by the bandwidth ($\approx t$), and therefore what results is still QPI with very clear and visible modulations.
@@
It is remarkable that the spectra remains homogeneous at both low and high energies even in the presence of a highly inhomogeneous order-parameter background. The reason for this in the $l<\xi$ case is that the quasiparticle response will be washed out over a coherence volume $\xi^{2}$, so any variations of the order parameter that are smaller than $\xi$ will be invisible in any observables that are directly derived from quasiparticle properties such as the LDOS. This can be seen again in Figs. 2 and 5: the extremely small-$\Delta_{T}$ regions do not exhibit a gapless spectrum, and instead show a finite $\Delta_{S}\approx\overline{\Delta}$. The superconducting quasiparticles in effect see a mostly homogeneous background with order parameter $\overline{\Delta}$ but are nevertheless scattered elastically by order-parameter inhomogeneity, which itself is weak on the scale set by the nearest-neighbor hopping amplitude $t$. The effect of this scattering will be to broaden the spectral functions at higher energies. This is corroborated by Fig. 3b: the LDOS at $\omega\approx 0.16$ is more broadly distributed than at lower energies, and consequently its average LDOS is lower than it would be in the absence of inhomogeneity. This is similar to what weak-impurity scattering in the Born limit predicts Lee-Hone et al. (2018), and suggests that the physics at $l<\xi$ is determined much more strongly by Born scattering (albeit in the “off-diagonal” channel) than in the other cases we will consider shortly, where the scattering is of a decidedly different nature.
@@ IV Intermediate-Sized Patches ($l\approx\xi$)
Figure 8: Plots of $\Delta_{T}$ (a), $\Delta_{S}$ (b), and $\Delta_{L}$ (c), and the absolute value of their respective Fourier transforms (d, e, and f) for patch size $l=8$. Only one disorder realization is presented. Shown here are quantities taken from the middlemost $256\times 256$ segment of the full $1024\times 256$ system.
@@
The real-space $\Delta_{T}$ map (Fig. 8a) shows that $\Delta_{T}$ is broadly distributed; regions with very small $\Delta_{T}$ coexist with regions where $\Delta_{T}$ is very large. The effect of this extended form of inhomogeneity in the order parameter in the spectral gap $\Delta_{S}$ (Fig. 8b) is striking: $\Delta_{S}$ has a broad distribution, like $\Delta_{T}$, but looks far splotchier than $\Delta_{T}$; the sharp square-shaped features seen in $\Delta_{T}$ are not present in these plots. Instead, $\Delta_{S}$ looks like a coarse-grained version of $\Delta_{T}$. It can be seen by eyeballing these two plots side by side that clusters of square patches where $\Delta_{T}$ is small show up in the $\Delta_{S}$ map as a spatially extended cluster where the spectra gap is small. An analogous observation can be made for the large-$\Delta_{T}$ clusters, but it can be seen that these show up in $\Delta_{S}$ as large-gap clusters that look smaller in spatial extent compared to the small-gap regions. While a clear correlation between $\Delta_{T}$ and $\Delta_{S}$ can be seen in these real-space maps, the correspondences are never exact between the two quantities. Importantly, it should be noted that while there exist regions where $\Delta_{T}\approx 0$, the spectral gap $\Delta_{S}$ never vanishes throughout the entire sample—the whole system exhibits a nonzero spectral gap, with the minimum spectral gap around $\approx 0.08$.
@@
$\Delta_{S}$ no longer exhibits the periodic modulations seen in the $l=1$ case. This serves to highlight a key difference between this case and the $l<\xi$ case considered earlier: the nature of the scattering of the Bogoliubov quasiparticles off of the order-parameter inhomogeneity is different for both of these cases. Because $l$ is now a few lattice sites long, only small-momentum scattering occurs. This means that order-parameter-induced QPI is a weaker effect for $l\approx\xi$. This can be seen explicitly by examining the Fourier transforms of $\Delta_{T}$ and $\Delta_{S}$ (Figs. 8d and e). The Fourier transform of $\Delta_{T}$ shows that strong intensities are localized around a relatively small portion of the Brillouin zone centered around $\mathbf{k}=0$, in contrast with what is seen for $l=1$ where the Fourier transform is featureless across the entire Brillouin zone; the resulting scattering off of the order-parameter inhomogeneity is primarily small-momentum and small-angle in nature. Higher harmonics are present in the Fourier transform of $\Delta_{T}$ owing to the uniformly square shape and size of the patches. The Fourier transform of $\Delta_{S}$ also shows similar characteristics: the intensity is largely localized near zero momentum, and the sharp, QPI-derived features seen in Fig. 2e for $l=1$ are no longer seen. However, the higher harmonics seen in the Fourier transform of $\Delta_{T}$ are no longer visible here as the patches in the $\Delta_{S}$ maps are more irregularly shaped.
@@
In contrast to the inhomogeneous nature of the $\Delta_{S}$ plots, the low-energy gap $\Delta_{L}$ (Fig. 2c) remains very much homogeneous, hardly deviating from $\Delta_{L}\approx 0.16$. Only faint traces of the large variation in $\Delta_{T}$ can be found in $\Delta_{L}$. Regions where many large-$\Delta_{T}$ patches are in close proximity to one another will give rise to a larger $\Delta_{L}$, and vice versa, but these effects are extremely faint and certainly much more muted than in the case of $\Delta_{S}$. The Fourier transform of $\Delta_{L}$ (Fig. 2f) features strong signals only at small wavevectors, in contrast to the $l=1$ case, indicating that for $l\approx\xi$, QPI even at low energies is affected by the presence of spatially extended order-parameter inhomogeneity, with the only real-space modulations present arising from small-momentum scattering processes. As $\Delta_{L}$ is a low-energy proxy for the superconducting order parameter and $\Delta_{S}$ a high-energy one, the difference in their behavior for $l=8$ stands in contrast to that seen in the $l=1$ case, for which both order-parameter proxies are highly homogeneous (with the only differences stemming from the differing nature of QPI at low and high energies). The dichotomy between $\Delta_{L}$ and $\Delta_{S}$ suggests a fundamental difference between the low-energy and high-energy states of an inhomogeneous $d$-wave superconductor when the inhomogeneity length scale $l\approx\xi$.
@@
At $\omega=0.16$ (Fig. 10c), we can see that the large-LDOS regions are no longer easily identified with any particular set of patches in Fig. 10b, and no longer retain the characteristic square shape of the underlying order-parameter patches, appearing more irregularly shaped instead. In addition to the large-LDOS regions, we identify two more categories of regions of interest here. The first are regions which correspond to a large local $\Delta_{S}$. Similar to the $\omega=0.09$ plot, the LDOS for these regions remains suppressed and is notably featureless, with no visible modulations within this regions. The second are the regions whose LDOS had peaked at energies $\omega<0.16$. At $\omega=0.16$, each of these regions consistently features a ring of enhanced intensity surrounding a region where the LDOS is suppresed. Finally, at $\omega=0.31$ (Fig. 10d), the remaining large-LDOS regions have a smaller value of the LDOS compared to analogous regions at smaller frequencies, consistent with the observation in Fig. 9a that the coherence-peak height is smaller for larger spectral gaps. These large-gap regions do not retain any sharp signatures of the square patches in $\Delta_{T}$. They appear instead as “islands” amid a background that hosts visible modulations. The large areas that host these modulations consist of all the regions whose local $\Delta_{S}$ is smaller than $\omega=0.31$; these modulations arise from quasiparticle scattering off of the large order-parameter “barriers” formed by the remnant large-$\Delta_{T}$ regions. At this energy, these modulations are visible because the spatial extent of the all regions whose local spectral gap $\Delta_{S}<0.31$ is large, on the order of several Fermi wavelengths; this is to be contrasted with the relative absence of modulations in the $\omega=0.16$ case, for which the regions with $\Delta_{S}<0.16$ are much smaller in extent.

### v2

@@ II.4 Order Parameter Proxies
A central motivation of this paper is figuring out the extent to which the various observables that can be extracted from $dI/dV$ measurements can serve as accurate proxies for the true underlying superconducting order parameter. In much of the STS literature, the energy at which the superconducting coherence peak is found—the “spectral gap”—has long served as a convenient proxy for the superconducting gap (on the other hand, it should be noted that many other STS papers interpret the spectral gap as the pseudogap energy). For a bulk superconductor in the clean limit, this correspondence is exact. However, in the presence of disorder, this is no longer automatically true. When chemical-potential disorder is present, the correlation between the spectral gap and the underlying order parameter is not strong, and in certain circumstances becomes very weak [44]. As this paper focuses purely on order-parameter inhomogeneity, we reexamine the correlation between the spectral gap and the order parameter in detail, and in particular focus on the role played by the inhomogeneity length scale $l$ on this correlation.
@@
We focus on two proxies for the local superconducting order parameter. The first one is simply the aforementioned spectral gap $\Delta_{S}$ and is obtained by finding the location of the positive-energy peak in the LDOS. (The negative-energy spectra are disregarded as these are contaminated by the presence of the Van Hove singularity right below the Fermi energy.) The second one is christened the “low-energy gap” $\Delta_{L}$ and is obtained by taking the slope of the positive-energy LDOS near zero energy. In a bulk clean $d$-wave superconductor, the slope near $\omega=0$ is inversely proportional to the bulk order parameter, so one can use the former to obtain the latter. We use this relationship $d\rho/d\omega|_{\omega\approx 0^{+}}\propto 1/\Delta_{L}$ to obtain an estimate for $\Delta_{L}$ from the low-energy slope of the LDOS at each lattice site. To be more precise, because we use a finite frequency grid in evaluating the LDOS, we use the finite difference quotient to approximate the low-energy slope of the LDOS:
@@
While this estimate is exact only for a homogeneous $d$-wave superconductor, we will nevertheless use Eqs. 10 and 11 to estimate the low-energy gap $\Delta_{L}$ for each LDOS spectrum for the inhomogeneous case.
These two proxies are to be compared with the “true” superconducting order parameter $\Delta_{T}$, which we define in the following way:
@@
We also retain information about the coherence-peak height $\rho_{C}(\mathbf{r}_{i})=\rho(\mathbf{r}_{i},\omega=+\Delta_{S,i})$ as its correlations with the order-parameter proxies can potentially provide valuable clues about the nature of the order-parameter inhomogeneity in real-world cuprates. In particular, it is known that $\rho_{C}$ and $\Delta_{S}$ are negatively correlated for many cuprates studied by STS such as Bi-2212 and Bi-2201 [5, 6, 8, 10, 51], so this provides an important experimental constraint on various theoretical models of order-parameter inhomogeneity.
@@ III Small Patches ($l<\xi$)
Figure 1: Plots of $\Delta_{T}$ (a), $\Delta_{S}$ (c), and $\Delta_{L}$ (e), and the absolute value of their respective Fourier transforms (b, d, and f) for patch size $l=1$. Only one disorder realization is presented. Shown here are quantities taken from the middlemost $256\times 256$ segment of the full $1024\times 256$ system.
@@
We first turn our attention to small patches characterized by $l<\xi$. To exemplify this, we show in Fig. 1 real-space and Fourier-transformed plots of $\Delta_{T}$, $\Delta_{S}$, and $\Delta_{L}$ for a $256\times 256$ subset of the system for the case where the mean order parameter is $\overline{\Delta}=0.16$ and the patch size is $l=1$. As mentioned earlier, for $\overline{\Delta}=0.16$, the coherence length is $\xi\approx 4.53$, so the patches are parametrically smaller than the coherence length. It can be seen from the real-space plot of the order parameter $\Delta_{T}$ (Fig. 1a) that it has a broad distribution ranging from $\Delta_{T}\approx 0$ to $\Delta_{T}\approx 0.3$, and that its modulations are short-ranged, with $\Delta_{T}$ on nearby sites largely uncorrelated with one another, except for rare regions where multiple patches with similar values are close to one another. The distribution of $\Delta_{T}$ is broad in that both regions with very small gaps and regions with very large gaps are present, often in close proximity with one another. In contrast, the spectral gap $\Delta_{S}$ (Fig. 1c) is comparatively narrowly distributed around $0.16$, except for a very small number of regions whose value clusters around $0.25$. The latter effect is due to the Van Hove singularity (VHS); in the clean limit with $\overline{\Delta}=0.16$, the VHS gives rise to a small shoulder at $\omega\approx 0.25$ above the positive-energy coherence peak at $\omega=0.16$. In the inhomogeneous setting, it appears that interference effects can cause this shoulder to mix with the coherence peak, leading to the formation of a peak where the shoulder should be. These clusters tend to form at large-$\Delta_{T}$ regions, but this effect appears to be rare as only a small number of large-$\Delta_{T}$ regions exhibit this effect at all.
@@
The real-space plot of $\Delta_{S}$ is characterized by weak modulations with a wavelength of around a few lattice sites. If one closely examines the $\Delta_{S}$ and $\Delta_{T}$ plots side by side, it can be seen that the two maps do not look similar to one another at all. While some large-$\Delta_{T}$ areas coincide with areas where $\Delta_{S}$ is large, the periodic modulations seen in $\Delta_{S}$ do not at all correspond to any modulations seen in $\Delta_{T}$. These periodic modulations are instead due to QPI arising from the presence of short-range disorder in the pairing (or “off-diagonal”) channel. The Fourier transform of the $\Delta_{S}$ map (Fig. 1d) confirms this: the dominant wavevectors here are precisely the scattering wavevectors responsible for QPI near the $d$-wave gap edge. This explanation is not altogether surprising. In the $l<\xi$ regime, the order-parameter inhomogeneity can be regarded as being in the pointlike limit in which the allowed scattering processes are unrestricted except for the requirement that they connect quasiparticle states for which the order parameter has the same sign. In the $l=1$ case, the modulations in the LDOS at $\omega\approx 0.16$ themselves form the modulations in the $\Delta_{S}$ maps (recall that $\Delta_{S}$ is the location of the positive-energy coherence peak). A close examination of Fig. 1d shows that the dominant scattering wavevectors giving rise to the modulations in the $\Delta_{S}$ map are the so-called sign-preserving $\mathbf{q}$-vectors” well-known from the octet-model description of QPI: $\mathbf{q}_{1}$, $\mathbf{q}_{4}$, and $\mathbf{q}_{5}$ [40, 63, 64].
@@
$\Delta_{L}$ (Fig. 1e) meanwhile can be seen to be very narrowly distributed around 0.16 as well, and does not reflect the comparatively large spatial variations in the underlying order parameter $\Delta_{T}$. Modulations can be seen here as well, but these are evidently very weak, and have a different characteristic length scale that those seen in $\Delta_{S}$. The Fourier transform of $\Delta_{L}$ (Fig. 1f) shows that these modulations arise from QPI as well, but this time at low energies near $\omega\approx 0$. The Fourier transforms (with visible QPI scattering wavevectors) of the two order-parameter proxies in Figs. 1d and 1f are in striking contrast to the Fourier transform of the true order parameter $\Delta_{T}$ (Fig. 1b), which is by and large featureless.
@@
Fig. 2b shows a two-dimensional histogram of the LDOS values across the same ensemble of spectra shown in the previous plot, with the disorder- and system-averaged LDOS shown in white. For each value of the frequency $\omega$, the distribution of the LDOS is obtained, and the density of each possible LDOS value is shown in Fig. 2b. It can be seen that the LDOS distribution does not deviate much from the average LDOS. This is especially true at low energies. As the energy is increased, one can see that the distribution becomes broader as $\to\pm\overline{\Delta}$, but the values remain within a narrow region surrounding the average LDOS.
More insight into the homogeneous nature of the states comes by way of plotting the LDOS itself (Fig. 3) taken from the same field of view as in Fig. 1 at four representative energies, which we take to be: $\omega=0.03$ (Fig. 3a), to represent the low-energy LDOS; $\omega=0.15$ and $\omega=0.17$ (Figs. 3b and d), which are the spectral gaps of the LDOS averaged over the first and last bins, respectively, in Fig. 2b; and $\omega=0.16$, which corresponds to the average superconducting order parameter in the system. The latter three energies are chosen to exemplify the high-energy states where inhomogeneity is strongest. The low-energy LDOS (Fig. 3a) is very featureless and does not contain any strong variations reflective of the underlying order parameter. As for the high-energy plots (Figs. 3b-d), these all are very similar to one another in that they feature modulations due to QPI. These modulations are particularly pronounced for $\omega=0.15$ (Fig. 3b), but they appear to weaken as energy is increased from this point. These LDOS patterns are consistent with QPI due to weak, narrowly distributed order-parameter inhomogeneity as previously studied by Sulangi et al. [64], but it is nevertheless striking that the same patterns are still present here even when the width of the order-parameter disorder distribution is parametrically much larger than in the case previously studied. This may be attributed to the fact that even with the relatively large width of the order-parameter inhomogeneity distribution we have used (on the order of the mean superconducting order parameter itself), it is still weak on the overall scale set by the bandwidth ($\approx t$), and therefore what results is still QPI with very clear and visible modulations.
@@
It is remarkable that the spectra remains homogeneous at both low and high energies even in the presence of a highly inhomogeneous order-parameter background. The reason for this in the $l<\xi$ case is that the quasiparticle response will be washed out over a coherence volume $\xi^{2}$, so any variations of the order parameter that are smaller than $\xi$ will be invisible in any observables that are directly derived from quasiparticle properties such as the LDOS. This can be seen again in Figs. 1 and 4: the extremely small-$\Delta_{T}$ regions do not exhibit a gapless spectrum, and instead show a finite $\Delta_{S}\approx\overline{\Delta}$. The superconducting quasiparticles in effect see a mostly homogeneous background with order parameter $\overline{\Delta}$ but are nevertheless scattered elastically by order-parameter inhomogeneity, which itself is weak on the scale set by the nearest-neighbor hopping amplitude $t$. The effect of this scattering will be to broaden the spectral functions at higher energies. This is corroborated by Fig. 2b: the LDOS at $\omega\approx 0.16$ is more broadly distributed than at lower energies, and consequently its average LDOS is lower than it would be in the absence of inhomogeneity. This is similar to what weak-impurity scattering in the Born limit predicts [65], and suggests that the physics at $l<\xi$ is determined much more strongly by Born scattering (albeit in the “off-diagonal” channel) than in the other cases we will consider shortly, where the scattering is of a decidedly different nature.
@@ IV Intermediate-Sized Patches ($l\approx\xi$)
Figure 7: Plots of $\Delta_{T}$ (a), $\Delta_{S}$ (c), and $\Delta_{L}$ (e), and the absolute value of their respective Fourier transforms (b, d, and f) for patch size $l=8$. Only one disorder realization is presented. Shown here are quantities taken from the middlemost $256\times 256$ segment of the full $1024\times 256$ system.
@@
The real-space $\Delta_{T}$ map (Fig. 7a) shows that $\Delta_{T}$ is broadly distributed; regions with very small $\Delta_{T}$ coexist with regions where $\Delta_{T}$ is very large. The effect of this extended form of inhomogeneity in the order parameter in the spectral gap $\Delta_{S}$ (Fig. 7c) is striking: $\Delta_{S}$ has a broad distribution, like $\Delta_{T}$, but looks far more splotchy than $\Delta_{T}$; the sharp square-shaped features seen in $\Delta_{T}$ are not present in these plots. Instead, $\Delta_{S}$ looks like a coarse-grained version of $\Delta_{T}$. It can be seen by eyeballing these two plots side by side that clusters of square patches where $\Delta_{T}$ is small show up in the $\Delta_{S}$ map as a spatially extended cluster where the spectra gap is small. An analogous observation can be made for the large-$\Delta_{T}$ clusters, but it can be seen that these show up in $\Delta_{S}$ as large-gap clusters that look smaller in spatial extent compared to the small-gap regions. While a clear correlation between $\Delta_{T}$ and $\Delta_{S}$ can be seen in these real-space maps, the correspondences are never exact between the two quantities. Importantly, it should be noted that while there exist regions where $\Delta_{T}\approx 0$, the spectral gap $\Delta_{S}$ never vanishes throughout the entire sample—the whole system exhibits a nonzero spectral gap, with the minimum spectral gap around $\approx 0.08$.
@@
$\Delta_{S}$ no longer exhibits the periodic modulations seen in the $l=1$ case. This serves to highlight a key difference between this case and the $l<\xi$ case considered earlier: the nature of the scattering of the Bogoliubov quasiparticles off of the order-parameter inhomogeneity is different for both of these cases. Because $l$ is now a few lattice sites long, only small-momentum scattering occurs. This means that order-parameter-induced QPI is a weaker effect for $l\approx\xi$. This can be seen explicitly by examining the Fourier transforms of $\Delta_{T}$ and $\Delta_{S}$ (Figs. 7b and d). The Fourier transform of $\Delta_{T}$ shows that strong intensities are localized around a relatively small portion of the Brillouin zone centered around $\mathbf{k}=0$, in contrast with what is seen for $l=1$ where the Fourier transform is featureless across the entire Brillouin zone; the resulting scattering off of the order-parameter inhomogeneity is primarily small-momentum and small-angle in nature. Higher harmonics are present in the Fourier transform of $\Delta_{T}$ owing to the uniformly square shape and size of the patches. The Fourier transform of $\Delta_{S}$ also shows similar characteristics: the intensity is largely localized near zero momentum, and the sharp, QPI-derived features seen in Fig. 1d for $l=1$ are no longer seen. However, the higher harmonics seen in the Fourier transform of $\Delta_{T}$ are no longer visible here as the patches in the $\Delta_{S}$ maps are more irregularly shaped.
@@
In contrast to the inhomogeneous nature of the $\Delta_{S}$ plots, the low-energy gap $\Delta_{L}$ (Fig. 1e) remains very much homogeneous, hardly deviating from $\Delta_{L}\approx 0.16$. Only faint traces of the large variation in $\Delta_{T}$ can be found in $\Delta_{L}$. Regions where many large-$\Delta_{T}$ patches are in close proximity to one another will give rise to a larger $\Delta_{L}$, and vice versa, but these effects are extremely faint and certainly much more muted than in the case of $\Delta_{S}$. The Fourier transform of $\Delta_{L}$ (Fig. 1f) features strong signals only at small wavevectors, in contrast to the $l=1$ case, indicating that for $l\approx\xi$, QPI even at low energies is affected by the presence of spatially extended order-parameter inhomogeneity, with the only real-space modulations present arising from small-momentum scattering processes. As $\Delta_{L}$ is a low-energy proxy for the superconducting order parameter and $\Delta_{S}$ a high-energy one, the difference in their behavior for $l=8$ stands in contrast to that seen in the $l=1$ case, for which both order-parameter proxies are highly homogeneous (with the only differences stemming from the differing nature of QPI at low and high energies). The dichotomy between $\Delta_{L}$ and $\Delta_{S}$ suggests a fundamental difference between the low-energy and high-energy states of an inhomogeneous $d$-wave superconductor when the inhomogeneity length scale $l\approx\xi$.
@@
At $\omega=0.16$ (Fig. 9c), we can see that the large-LDOS regions are no longer easily identified with any particular set of patches in Fig. 9b, and no longer retain the characteristic square shape of the underlying order-parameter patches, appearing more irregularly shaped instead. In addition to the large-LDOS regions, we identify two more categories of regions of interest here. The first are regions which correspond to a large local $\Delta_{S}$. Similar to the $\omega=0.09$ plot, the LDOS for these regions remains suppressed and is notably featureless, with no visible modulations within this regions. The second are the regions whose LDOS had peaked at energies $\omega<0.16$. At $\omega=0.16$, each of these regions consistently features a ring of enhanced intensity surrounding a region where the LDOS is suppressed. Finally, at $\omega=0.31$ (Fig. 9d), the remaining large-LDOS regions have a smaller value of the LDOS compared to analogous regions at smaller frequencies, consistent with the observation in Fig. 8a that the coherence-peak height is smaller for larger spectral gaps. These large-gap regions do not retain any sharp signatures of the square patches in $\Delta_{T}$. They appear instead as “islands” amid a background that hosts visible modulations. The large areas that host these modulations consist of all the regions whose local $\Delta_{S}$ is smaller than $\omega=0.31$; these modulations arise from quasiparticle scattering off of the large order-parameter “barriers” formed by the remnant large-$\Delta_{T}$ regions. At this energy, these modulations are visible because the spatial extent of the all regions whose local spectral gap $\Delta_{S}<0.31$ is large, on the order of several Fermi wavelengths; this is to be contrasted with the relative absence of modulations in the $\omega=0.16$ case, for which the regions with $\Delta_{S}<0.16$ are much smaller in extent.